OpenStudy (automaticloveletter):

Two equations are given below: m + 3n = 10 m = n − 2 What is the solution to the set of equations in the form (m, n)? (1, 3) (2, 4) (0, 2) (4, 6)

OpenStudy (chillout):

We have \(m=n-2\). Just plug that in the first equation and we will have \((n-2)+3n=10\) which is in turn \(4n=12\). We have that \(n=3\) and \(m=3-2=1\)

OpenStudy (chillout):

the solution in the form of (m,n) is given by (1,3)

OpenStudy (automaticloveletter):

Can you check my last question?

OpenStudy (chillout):

Sure. Just link it to me :)